OpenStudy (anonymous):

Combine like terms. 8 + 16r – 9r A. 8 + 7r B. 15r C. 15 D. 8 – 7r

OpenStudy (anonymous):

@AskMeLotss

OpenStudy (anonymous):

just a sec.

OpenStudy (alexandervonhumboldt2):

8+16r-9r=8+(16-9)r=8+7r

OpenStudy (anonymous):

8 + 7r this is the answer

OpenStudy (anonymous):

anymore questions?

OpenStudy (anonymous):

yay